Using the values, X = 0, 1, 2, 4, 6 and respectively P(X) = .1, .2, ?, .3, .3 find the probability that X is greater than 1.

A) .6
B) .1
C) .7
D) .3

Answer by Edwin McCravy(20064)

X    P(X)
--------- 
0    .1
1    .2 
2     ?
4    .3
6    .3

For this to be a probability distribution, the sum
of the probabilities must = 1.  So since the probability
of the other four values of X have sum
.1+.2+.3+.3 = .9, then P(2) = 1.0-.9 = .1 and the
distribution is:

X    P(X)
--------- 
0    .1
1    .2 
2    .1
4    .3
6    .3
--------
Sum = 1

So P(X > 1) = P(X=2)+P(X=4)+P(X=6) = .1+.3+.3 = .7,

Or what would have been the same, 

P(X > 1) = 1 - P(X < 2) = 1-P(X=0)-P(X=1) = 1-.1-.2 = .3

Edwin